OpenStudy (anonymous):

Find an equation of the line passing through the point (6,5) and perpendicular to the line whose equation is 2y + 3x = 6. what is it please i need help

OpenStudy (amistre64):

swap your x and y numbers and negate one of them 3y -2x = # ; now fill in the point to get a new sum

OpenStudy (amistre64):

3(5) -2(6) = # 15 -12 = # 3 = # .................. 3y -2x = 3 should be the equation then

OpenStudy (anonymous):

thanxz

OpenStudy (amistre64):

yep :)